A trolley of mass 200 kg moves with a unifo s1 rm speed of 36 km/h on a frictionless track. speed A child of mass 20 kg runs on the trolley from one end to the other (10 m away) with a of 4 m relative to the trolley in a direction opposite to the its motion, and the trolley. i. What is th ii. e final speed of the trolley ? How much has the trolley moved from t jumps out of he time the child begins to run
1
Expert's answer
2019-10-21T08:54:58-0400

From the conservation of momentum:


"(m+M)V=m(v-u)+Mv"

"(20+200)\frac{36}{3.6}=20(v-4)+200v"

The final speed of the trolley:

"v=10.4\frac{m}{s}"

"t=\frac{d}{v}=\frac{10}{4}=2.5\ s"

Thus,


"s=v_1t=(10.4)(2.5)=26\ m"